OpenStudy (anonymous):

the ratio of boys to girls in Mrs.Wilson's class is 4 to 3. If there are 28 students in a class, how many boys are in the class?

OpenStudy (luigi0210):

I think it might be 16

hero (hero):

4 + 3 = 7 4/7 = x/28 x = 16